OpenStudy (anonymous):

question on this problem.....m-8n+m+8n=

OpenStudy (anonymous):

2m

OpenStudy (cwrw238):

you can only add/subtract like terms if you bring them together it will help m + m - 8n + 8n =

OpenStudy (anonymous):

im still learning and my instructor teaches to where i dont understand

OpenStudy (cwrw238):

m + m = 2m and -8n + 8n = 0 o answer is 2m

OpenStudy (cwrw238):

does that help you to understand?

OpenStudy (anonymous):

yes

OpenStudy (cwrw238):

great